# Planning-Driven Programming: A Large Language Model Programming Workflow

### v1

@@ 4 LPW with Sampling
Baseline 74.4 - 66.5 - 67.4 - 52.8 -
SP 77.4 3.1 69.5 3.1 69.2 1.8 52.4 -0.4
SD 81.1 6.7 72.0 5.5 71.2 3.8 56.0 3.2
LDB 82.9 8.5 72.6 6.1 72.4 5.0 55.6 2.8
LPW (ours) 89.0 14.6 77.4 11.0 76.0 8.6 57.6 4.8
GPT-3.5 SLPW (ours) 89.6 15.2 77.4 11.0 77.2 9.8 58.2 5.4
Baseline 73.2 - 61.0 - 44.0 - 35.4 -
SP 78.0 4.9 65.2 4.3 48.6 4.6 38.4 3.0
SD 81.7 8.5 68.3 7.3 63.6 19.6 50.0 14.6
LDB 84.1 11.0 72.0 11.0 57.2 13.2 44.8 9.4
LPW (ours) 88.4 15.2 76.2 15.2 73.6 29.6 56.4 21.0
Llama-3 SLPW (ours) 89.0 15.9 76.2 15.2 75.0 31.0 57.2 21.8
Baseline 36.0 - 32.3 - 39.0 - 33.2 -
SP 40.9 4.9 34.8 2.4 46.4 7.4 37.6 4.4
SD 51.2 15.2 45.7 13.4 45.8 6.8 36.6 3.4
LDB 65.9 29.9 54.9 22.6 52.4 13.4 42.8 9.6
LPW (ours) 76.8 40.9 62.8 30.5 64.0 25.0 48.4 15.2
Phi-3 SLPW(ours) 81.1 45.1 67.1 34.8 69.6 30.6 52.2 19.0
@@
Table 1: Comparisons of Baseline, Self-Planning (SP), Self-Debugging (+Expl) (SD), LDB, LPW and SLPW in terms of Pass@1 accuracy (Acc) and improvement ($\Delta$) with respect to Baseline across benchmarks HumanEval, HumanEval-ET, MBPP, and MBPP-ET with LLMs GPT-3.5, Llama-3, and Phi-3. Acc and $\Delta$ are measured in percentages. Results for LPW and SLPW are in bold, and the best results are highlighted in red.
@@ 5 Experiments
Benchmarks. We evaluate LPW and SLPW on the well-established text-to-code benchmarks HumanEval, MBPP, HumanEval-ET, and MBPP-ET, where the given context outlines the intended functionality of the program to be synthesized. HumanEval-ET and MBPP-ET introduce approximately 100 additional hidden tests, covering numerous edge cases for each problem in HumanEval and MBPP, thus being regarded as more reliable benchmarks for code evaluation  (Dong et al., 2023a). In HumanEval and HumanEval-ET, we treat the test cases described in the task description as visible tests, typically 2-5 per task. For MBPP, we consider its test set that contains 500 problems with 3 hidden tests per problem. We set the first hidden test as the visible test and treat the other two as hidden, consistent with studies (Chen et al., 2023b; Zhong et al., 2024; Ni et al., 2023; Shi et al., 2022). MBPP-ET uses the same set of problems and visible tests for each problem as MBPP.
@@
Experimental Setup. We compare LPW and SLPW with the representative code generation approaches Self-Planning (SP) (Jiang et al., 2023) , Self-Debugging (+Expl) (SD)  (Chen et al., 2023b) and Large Language Model Debugger (LDB) (Zhong et al., 2024). SP relies solely on the LLM-generated solution plan to produce the program solution in a single attempt without refinements. SD uses a rubber duck debugging approach in LLMs, where LLMs are prompted to provide explanations of generated programs as feedback for debugging. LDB, a state-of-the-art LLM debugger, segments generated programs into blocks based on the control flow graph, which facilitates bug detection and the refinement of each program block using runtime execution information in LLMs. A detailed comparison between the baseline methods and our methods are summarized in Tables 14 and 15 in the Appendix. We generate the seed programs with the same prompts and parameters introduced by Chen et al. (2023b) for SD and LDB and label the performance of seed programs as Baseline. We note that SD and LDB only perform refinements on seed programs that fail the visible tests. We experiment with various LLMs with different parameter sizes, including GPT-3.5 (turbo-0125, $\geq$175B), Llama-3 (70B-Instruct), and Phi-3 (14B-Instruct) to evaluate performance and demonstrate that both LPW and SLPW are model-independent.
We use the Pass@1 accuracy as the evaluation metric. We apply 2-shot prompting in both LPW and SLPW, with maximum 12 iterations for the solution generation phase and the code implementation phase, respectively. Similarly, we set the maximum number of debugging iterations to 12 for SD and LDB. In SLPW, $k$ is configured as 6, and $q$ is set to 3. For instance, the solution generation phase initially produces 6 plan samples. Subsequently, first 3 solution plans along with their verifications are returned within 12 iterations, or $[0,3)$ solution plans and their verifications are provided as the output upon completing 12 iterations. All following experiments adhere to these parameter settings unless otherwise specified. Empirical discussion on parameters is available in Appendix A.2.
@@
Main Results. Table 1 presents the Pass@1 accuracy for Baseline, SP, SD, LDB, LPW, and SLPW, along with their respective improvements over Baseline. Overall, LPW and SLPW consistently outperform all competing methods across all benchmarks and with various LLM backbones, showcasing the effectiveness of the proposed workflow and demonstrating the model-independent benefit of LPW and SLPW. Compared to LDB, LPW improves Pass@1 accuracy by 6.1%, 4.9% 3.6%, and 2%, on HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively, with the GPT-3.5 backbone and achieves up to 16.4% improvement on MBPP when using Llama-3 as the backbone. LPW achieves the same performance as SLPW on HumanEval-ET when leveraging GPT-3.5 and Llama-3 as the backbones. SLPW slightly surpasses LPW by around 1% across all benchmarks, and achieves the best accuracy: 89.6% for HumanEval, 77.4% for HumanEval-ET, 77.2% for MBPP and 58.2% for MBPP-ET with GPT-3.5. Moreover, when using Phi-3 as the backbone, SLPW shows the highest improvement up to 5.6% over LPW on MBPP and up to 45.1% over Baseline on HumanEval. Compared with HumanEval and MBPP, all approaches perform worse on HumanEval-ET and MBPP-ET across different LLM backbones as thorough edge cases are contained in the hidden tests. This result is consistent with previous work (Dong et al., 2023b; Lin et al., 2024; Mu et al., 2023). The detailed failure analysis is available in the Appendix A.4.
@@
Baseline 91.5 81.7 78.4 62.6 41.7 28.0
LDB 92.1 81.7 82.4 65.4 53.2 29.3
GPT-4o LPW (ours) 97.0 84.1 84.8 65.8 62.6 34.7
SLPW (ours) 98.2 84.8 84.8 66.0 64.0 35.3
@@
Table 2: Pass@1 accuracy for Baseline, LDB, LPW, and SLPW on the same benchmarks in Table 1, as well as APPS and CodeContests when using the LLM GPT-4o (2024-05-13) as the backbone.
Results on Advanced LLM with Competitive Benchmarks. To further demonstrate the effectiveness of LPW and SLPW, we evaluate their performance against LDB on the same benchmarks presented in Table 1, as well as on two competitive benchmarks, APPS and CodeContests, using the advanced LLM GPT-4o as the backbone. For APPS and CodeContests, we use subsets of 139 and 150 problems, respectively. APPS and CodeContests are unstructured benchmarks where visible tests are intermingled with the problem statements and function signatures are excluded. To align input data structure across benchmarks, we instruct GPT-4o to derive the optimal function signature and identify visible tests for each problem prior to conducting experiments. The experiment results are shown in Table 2 and the Pass@1 accuracy is reported. Similarly, the performance of the seed programs for LDB is referred to as Baseline. LPW outperforms Baseline and LDB across all benchmarks, achieving the same 84.8% accuracy as SLPW on MBPP. SLPW further improves performance and establishes new state-of-the-art Pass@1 accuracy across all benchmarks, notably achieving 98.2% on HumanEval. The outstanding performance of SLPW indicates that sampling and debugging are mutually complementary in enhancing code generation performance. For APPS and CodeContests, LPW and SLPW achieve over 62% and 34% accuracy, respectively, surpassing LDB by around 10% and 5% accuracy, highlighting the advantages of LPW and SLPW in tackling challenging benchmarks. GPT-4o is considered as a more powerful LLM. Baseline achieves 91.5% and 28% accuracy on HumanEval and CodeContests without debugging, while LDB shows a negligible improvement of only 0.6% and 1.3% compared to Baseline on these two benchmarks. This underscores the limitations of debugging with coarse feedback. In contrast, the intended solution with respective to visible tests represented by the plan verification allows LPW and SLPW to clarify issues before code generation and efficiently correct bugs overlooked by LLMs. Appendix A.7 discusses problems GPT-4o fails to address and structured examples from APPS and CodeContests.
@@
Table 3: The impact on Pass@1 accuracy with additional visible tests using the GPT-3.5 backbone. MBPP-ET-3 includes two more visible tests per problem than MBPP-ET. $\Delta$ represents the accuracy improvement of MBPP-ET-3 over MBPP-ET. Pass@1 accuracy and $\Delta$ are measured as percentages.
Learning from Test. We further investigate the impact of the number of visible tests on SD, LDB, LPW, and SLPW that use visible tests to refine code. We propose a variant of MBPP-ET, denoted as MBPP-ET-3. In MBPP-ET-3, each problem’s visible tests are the three hidden tests from MBPP, while the hidden tests are the extended test cases introduced in MBPP-ET. In other words, each problem in MBPP-ET-3 contains two more visible tests than in MBPP-ET, thereby providing informative feedback for better bug identification and program refinement in LLMs. Table 3 compares the Pass@1 accuracy of SD, LDB, LPW and SLPW on the MBPP-ET-3 benchmark with the GPT-3.5 backbone. LPW and SLPW dominate in both accuracy and improvement. SLPW achieves the highest accuracy of 63.4% on MBPP-ET-3 and the largest improvement of 5.2% over MBPP-ET. LPW and SLPW exploit visible tests by producing the step-by-step solutions for each visible test to clarify initial code logic and inform subsequent refinements, demonstrating superior efficiency in utilizing visible tests among the evaluated methods.
@@
Figure 4: The impact on Pass@1 accuracy with the increased number of code implementation iterations/debugging iterations (part (a)), and across different difficulty levels, Easy, Medium, and Hard (part (b)) on the HumanEval benchmark when leveraging GPT-3.5 as the LLM backbone.
@@
Performance Analysis. Figure 4 evaluates the Pass@1 accuracy of LPW and SLPW when considering different numbers of code implementation iterations (part (a)) and varying difficulty levels (part (b)) on the HumanEval benchmark when using GPT-3.5 as the backbone. For SD and LDB, we allocate the same number of debugging iterations. We note that all evaluated approaches start from iteration 0, representing the Pass@1 accuracy before debugging. Specifically, for SD and LDB, this reflects the seed program (Baseline) accuracy, while for LPW and SLPW, it indicates the accuracy after generating a program for each plan and its verification produced from the solution generation phase. In part (a) of Figure 4, Baseline and SP are plotted as straight lines with 74.4% and 77.4% accuracy, respectively, due to no debugging involved. Baseline and SP serve as the control group to illustrate when debugging methods surpass no-debugging methods. SD and LDB refine incorrect programs in Baseline, surpassing SP after two iterations. LPW starts debugging from an initial 79.9% accuracy, while SLPW begins from 84.8%. Both are higher than the 77.4% for SP, highlighting the importance of plan verification in initial code generation. LPW surpasses the best performance of SD and LDB after only one iteration, demonstrating its efficient code refinement strategy. The initial debugging accuracy of SLPW, 84.8%, exceeds the best performance of SD and LDB, showcasing the advantages of incorporating sampling. LPW and SLPW gradually refine the code and reach the highest Pass@1 accuracy by the 10th iteration. We apply the method described in Zhong et al. (2024) to generate the difficulty annotation in part (b) of Figure 4 by querying GPT-3.5 with problem descriptions and canonical solutions. LPW and SLPW display convincing performance, exceeding 85% accuracy across all difficulty levels. For the Hard level, LPW and SLPW achieve 85.7% and 90.5% accuracy, in contrast to competing approaches whose performance notably degrades to below 70%. This aligns with LPW and SLPW usage scenarios, where the precise natural language solution described in the plan and plan verification is essential for logical completeness and understanding non-trivial bugs in the program, particularly when problems involve complex logical reasoning steps. See Appendix A.5 and A.6 for a discussion on the accuracy of solution plans, plan verifications and refined programs generated by GPT-3.5.
@@
Table 4: Pass@1 accuracy (Acc) for different variants of LPW and SLPW with GPT-3.5. The suffixes -S and -C represent excluding the solution generation phase and the code implementation phase. $\Delta$ denotes the decrease against LPW and SLPW. Acc and $\Delta$ are measured in percentages.
Ablation Study. Table 4 shows the Pass@1 accuracy of different variants of LPW and SLPW on the HumanEval and MBPP benchmarks with GPT-3.5. The suffix -S stands for LPW and SLPW variants that exclude the solution generation phase, while the suffix -C represents the removal of the code implementation phase, specifically omitting code refinements. For each problem, LPW-S repairs the seed program that fails visible tests from Baseline, leveraging code explanations and runtime information but without plan and plan verification. LPW-C generates the program solution based on the plan and its verification without refinements. SLPW-S and SLPW-C maintain the same $k=6$ and $q=3$ settings as SLPW. SLPW-S follows the same refinement approach as LPW-S to repair $q$ seed programs, generated in the same way as Baseline. SLPW-C employs the same strategy as LPW-C to create a program for each plan and its verification, but without applying refinements. The results in Table 4 show that removing either phase in LPW and SLPW decreases their performance, indicating that both the solution generation phase and the code implementation phase are crucial for optimal performance. For MBPP, both phases exhibit a similar impact in LPW and SLPW. In contrast, LPW-C experiences a significant 9.1% decrease on the HumanEval benchmark compared to LPW, as debugging plays a crucial role in reaching the accurate solution given the large number of visible tests in HumanEval. This also underscores the benefit of debugging for maintaining code quality when sampling is omitted. Meanwhile, the smaller 4.8% decrease for SLPW-C compared to SLPW on HumanEval demonstrates the advantage of sampling in the absence of debugging. See Appendix A.3 for additional ablation study.

### v2

@@ 3 Workflow Structure
Baseline 74.4 ±$\pm 0.8$ 66.5 ±$\pm 1.3$ 67.4 ±$\pm 0.5$ 52.8 ±$\pm 0.3$
SP 77.4 ±$\pm 0.8$ 3.0 69.5 ±$\pm 0.8$ 3.0 69.2 ±$\pm 0.4$ 1.8 52.4 ±$\pm 0.2$0.4
MapCoder 77.4 ±$\pm 1.0$ 3.0 66.5 ±$\pm 1.0$ 0.0 72.0 ±$\pm 0.5$ 4.6 56.6 ±$\pm 0.5$ 3.8
SD 81.1 ±$\pm 1.0$ 6.7 72.0 ±$\pm 1.0$ 5.5 71.2 ±$\pm 0.3$ 3.8 56.0 ±$\pm 0.1$ 3.2
LDB 82.9 ±$\pm 1.0$ 8.5 72.6 ±$\pm 1.0$ 6.1 72.4 ±$\pm 0.3$ 5.0 55.6 ±$\pm 0.2$ 2.8
GPT-3.5 LPW (ours) 89.0 ±$\pm 0.8$ 14.6 77.4 ±$\pm 0.8$ 10.9 76.0 ±$\pm 0.2$ 8.6 57.6 ±$\pm 0.1$ 4.8
Baseline 73.2 ±$\pm 1.3$ 61.0 ±$\pm 1.0$ 44.0 ±$\pm 1.2$ 35.4 ±$\pm 1.0$
SP 78.0 ±$\pm 2.0$ 4.8 65.2 ±$\pm 1.0$ 4.2 48.6 ±$\pm 1.4$ 4.6 38.4 ±$\pm 1.4$ 3.0
MapCoder 83.5 ±$\pm 1.3$ 10.3 73.8 ±$\pm 0.8$ 12.8 71.4 ±$\pm 1.0$ 27.4 55.6 ±$\pm 1.0$ 20.2
SD 81.7 ±$\pm 1.3$ 8.5 68.3 ±$\pm 0.8$ 7.3 63.6 ±$\pm 1.2$ 19.6 50.0 ±$\pm 1.3$ 14.6
LDB 84.1 ±$\pm 1.7$ 10.9 72.0 ±$\pm 0.8$ 11.0 57.2 ±$\pm 1.6$ 13.2 44.8 ±$\pm 1.4$ 9.4
Llama-3 LPW (ours) 88.4 ±$\pm 1.6$ 15.2 76.2 ±$\pm 1.3$ 15.2 73.6 ±$\pm 1.3$ 29.6 56.4 ±$\pm 1.2$ 21.0
Baseline 36.0 ±$\pm 1.0$ 32.3 ±$\pm 1.0$ 39.0 ±$\pm 1.3$ 33.2 ±$\pm 1.4$
SP 40.8 ±$\pm 1.4$ 4.8 34.8 ±$\pm 0.9$ 2.5 46.4 ±$\pm 1.4$ 7.4 37.6 ±$\pm 1.4$ 4.4
MapCoder
SD 51.2 ±$\pm 1.2$ 15.2 45.7 ±$\pm 1.0$ 13.4 45.8 ±$\pm 1.2$ 6.8 36.6 ±$\pm 1.2$ 3.4
LDB 65.9 ±$\pm 1.6$ 29.9 54.9 ±$\pm 0.9$ 22.6 52.4 ±$\pm 1.6$ 13.4 42.8 ±$\pm 1.4$ 9.6
Phi-3 LPW (ours) 76.8 ±$\pm 1.3$ 40.8 62.8 ±$\pm 1.3$ 30.5 64.0 ±$\pm 1.2$ 25.0 48.4 ±$\pm 1.2$ 15.2
@@
Table 1: Comparisons of Baseline, SP, MapCoder, SD, LDB, and LPW in terms of Pass@1 accuracy (Acc) and improvement ($\Delta$) with respect to Baseline across benchmarks HumanEval, HumanEval-ET, MBPP, and MBPP-ET with LLMs GPT-3.5, Llama-3, and Phi-3. Acc and $\Delta$ are measured in percentages. Best results are highlighted in red. The standard deviation ($\pm$) is calculated based on three runs and applies to other experiments when reported.
@@ 4 Experiments
Benchmarks. We first evaluate LPW on the well-established text-to-code benchmarks HumanEval, MBPP, HumanEval-ET, and MBPP-ET, where the given context outlines the intended functionality of the program to be synthesized. HumanEval-ET and MBPP-ET introduce approximately 100 additional hidden tests, covering numerous edge cases, for each problem in HumanEval and MBPP, thus being regarded as more reliable benchmarks  (Dong et al., 2023a). In HumanEval and HumanEval-ET, we treat the test cases described in the task description as visible tests, typically 2-5 per task. For MBPP, we consider its test set that contains 500 problems with 3 hidden tests per problem. We set the first hidden test as the visible test and treat the other two as hidden, consistent with studies (Chen et al., 2023b; Zhong et al., 2024; Ni et al., 2023; Shi et al., 2022). MBPP-ET uses the same set of problems and visible tests for each problem as MBPP.
@@
Experimental Setup. We compare LPW with the representative code generation approaches Self-Planning (SP) (Jiang et al., 2023), MapCoder  (Islam et al., 2024), Self-Debugging (+Expl) (SD)  (Chen et al., 2023b), and Large Language Model Debugger (LDB) (Zhong et al., 2024). SP relies solely on the LLM-generated solution plan to produce the program solution in a single effort without refinements. MapCoder, a multi-agent collaborative approach, generates multiple unverified plans and traverses them to produce and refine code based on the current plan. SD uses a rubber duck debugging approach in LLMs, where LLMs are prompted to provide explanations of generated programs as feedback for debugging. LDB, a state-of-the-art LLM debugger, segments generated programs into blocks based on the control flow graph, which facilitates bug detection and the refinement of each program block using runtime information in LLMs. A detailed comparison of different methods is summarized in Appendix Table 15.
We generate a seed program for each problem with the same prompts and parameters introduced by Chen et al. (2023b) for SD and LDB and label the performance of seed programs as Baseline. We note that SD and LDB only perform refinements on the seed program that fails the visible tests. We experiment with various LLMs with different parameter sizes, including GPT-3.5 (turbo-0125, $\geq$175B), Llama-3 (70B-Instruct), and Phi-3 (14B-Instruct) to evaluate performance and demonstrate that LPW is model-independent. The Pass@1 accuracy is reported. We apply 2-shot prompting in LPW, with a maximum of 12 iterations for both the solution generation and code implementation phases. Similarly, we set the maximum debugging iterations to 12 for SD and LDB. MapCoder generates 3 plans, each with up to 4 refinement iterations. All following experiments adhere to these iteration settings. An empirical discussion on the number of iterations for LPW is available in Appendix B.
@@
Results on various LLMs. Table 1 presents the Pass@1 accuracy for evaluated approaches, along with their respective improvements over Baseline. LPW outperforms all competing methods across all benchmarks and with various LLM backbones. Compared to LDB, LPW improves Pass@1 accuracy by 6.1%, 4.8% 3.6%, and 2%, on HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively, with GPT-3.5 and achieves up to 16.4% improvement on MBPP with Llama-3. These results showcase the effectiveness of the proposed workflow and demonstrate the model-independent benefits of LPW. MapCoder fails on Phi-3 as it requires strict XML-formatted responses, which pose a challenge for Phi-3. The failure analysis for LPW with GPT-3.5 is available in Appendix D.
@@
Baseline 91.5 ±$\pm 0.3$ 81.7 ±$\pm 0.3$ 78.4 ±$\pm 0.4$ 62.6 ±$\pm 0.2$ 45.7 ±$\pm 0.6$ 41.7±$\pm 0.9$ 28.0 ±$\pm 0.5$
GPT-4o LDB 92.1 ±$\pm 0.0$ 81.7 ±$\pm 0.0$ 82.4 ±$\pm 0.3$ 65.4 ±$\pm 0.0$ 54.3 ±$\pm 0.3$ 53.2 ±$\pm 0.3$ 29.3 ±$\pm 0.3$
LPW (ours) 98.2 ±$\pm 0.3$ 84.8 ±$\pm 0.3$ 84.8 ±$\pm 0.2$ 65.8 ±$\pm 0.1$ 59.3 ±$\pm 0.6$ 62.6 ±$\pm 0.3$ 34.7 ±$\pm 0.3$
@@
Table 2: Pass@1 accuracy, in percentages, for Baseline, LDB, and LPW on HumanEval, HumanEval-ET, MBPP, MBPP-ET, LiveCode, APPS and CodeContests benchmarks when using GPT-4o (2024-05-13) as the backbone.
Results on Advanced LLM. To further demonstrate the effectiveness of LPW, we evaluate its performance on the same benchmarks presented in Table 1, along with the contamination-free benchmark, LiveCode, and two competitive benchmarks, APPS and CodeContests, using the advanced LLM GPT-4o as the backbone. LDB is compared due to its second-highest performance in Table 1 using GPT-3.5. We sample 140 problems from LiveCode, published between November 2023 and September 2024, postdating GPT-4o’s cutoff date. For APPS and CodeContests, we use subsets of 139 and 150 problems, respectively. The experiment results are shown in Table 2. Similarly, the performance of the seed programs for LDB is referred to as Baseline. LPW outperforms Baseline and LDB across all benchmarks, and establishes new state-of-the-art Pass@1 accuracy, notably achieving 98.2% on HumanEval. GPT-4o exhibits reduced performance on LiveCode, while LPW reliably outperforms LDB by 5% accuracy. For APPS and CodeContests, LPW surpasses LDB by around 10% and 5% accuracy, highlighting the advantages of LPW in tackling challenging benchmarks. Incorporating the plan verification allows LPW to clarify issues before code generation and efficiently correct bugs overlooked by LLMs. In contrast, LDB shows a negligible improvement of only 0.6% and 1.3% compared to Baseline on HumanEval and CodeContests, underscoring the limitations of debugging with coarse feedback. See Appendix G for tasks that LPW fails to address with GPT-4o.
@@
Table 3: The impact on Pass@1 accuracy with additional visible tests using the GPT-3.5 backbone. MBPP-ET-3 includes two more visible tests per problem than MBPP-ET. $\Delta$ represents the accuracy improvement on MBPP-ET-3 over MBPP-ET. Pass@1 accuracy and $\Delta$ are measured as percentages.
@@
Figure 4: The impact on Pass@1 accuracy with the increased number of code implementation iterations/debugging iterations on the HumanEval benchmark when leveraging GPT-3.5 as the LLM backbone. The shaded area represents the standard deviation.
Learning from Test. We further investigate the impact of the number of visible tests on MapCoder, SD, LDB, and LPW that use visible tests to refine code. We propose a variant of MBPP-ET, denoted as MBPP-ET-3. In MBPP-ET-3, each problem’s visible tests are the three hidden tests from MBPP, while the hidden tests are the extended test cases introduced in MBPP-ET. In other words, each problem in MBPP-ET-3 contains two more visible tests than in MBPP-ET. Results in Table 3 show that LPW achieves the highest Pass@1 accuracy of 62.0% on MBPP-ET-3 and the largest improvement of 4.4% over MBPP-ET. LPW produces the step-by-step solution for each visible test to clarify code development logic and inform subsequent refinements, demonstrating superior efficiency in utilizing visible tests to improve performance among evaluated methods.
@@
Performance Analysis. Figure 4 evaluates the Pass@1 accuracy of LPW when considering different numbers of code implementation iterations on the HumanEval benchmark with GPT-3.5. For SD and LDB, we allocate the same number of debugging iterations. We omit MapCoder due to its distinct refinement strategy. We note that all evaluated approaches start from iteration 0, representing the Pass@1 accuracy before debugging. In Figure 4, Baseline and SP are plotted as straight lines with 74.4% and 77.4% accuracy, respectively, due to no debugging involved. Baseline and SP serve as the control group to illustrate when debugging methods surpass no-debugging methods. SD and LDB refine incorrect programs in Baseline, surpassing SP after two iterations. LPW starts debugging from an initial 79.9% accuracy, higher than the 77.4% for SP, underscoring the importance of plan verification in initial code generation. LPW surpasses the best performance of SD and LDB after only one iteration, demonstrating its efficient code refinement strategy. LPW gradually refines the code and reaches the highest accuracy by the 10th iteration.
@@
Table 4: Pass@1 accuracy (Acc) for different variants of LPW with GPT-3.5. $\Delta$ denotes the decrease against LPW. Acc and $\Delta$ are measured in percentages. See Appendix 6 for additional ablation study.
Figure 5: Pass@1 accuracy vs. average token cost per program for LDB and LPW on different benchmarks using GPT-4o as the LLM backbone. K is $10^{3}$. The standard deviation is too small to be visible.
Figure 6: Pass@1 accuracy as a function of token consumption for LDB, Repeated Sampling, and LPW on the APPS and CodeContests benchmarks with GPT-4o. The standard deviation is omitted for better illustration. The same illustration for the LiveCode benchmark is available in Appendix Figure 8.
Ablation Study. Table 4 shows the Pass@1 accuracy of different variants of LPW on the HumanEval and MBPP benchmarks with GPT-3.5. The suffix -V denotes the exclusion of plan verification in both solution generation and code implementation phases; -S stands for the LPW variant that excludes the solution generation phase; while -C represents the removal of the code implementation phase, specifically omitting code refinements. For each problem, LPW-V generates the initial program based on the unverified plan and repairs the program using only code explanation and runtime information. LPW-S refines the seed program from Baseline that fails visible tests, leveraging only code explanation and runtime information. LPW-C generates the program solution based on the plan and its verification without refinements.
The performance decline of LPW-V highlights the significance of plan verification, which serves as the intended solution for visible tests, improving the performance of LLMs in both initial code generation and subsequent refinements. LPW-V considers the unverified plan when drafting initial programs. However, the effect of the unverified plan is limited, as LPW-V shows only slight improvement on MBPP compared to LPW-S, which excludes both the plan and plan verification. This aligns with the results in Table 1, where Self-Planning shows minimal improvement compared to Baseline. The results of LPW-S and LPW-C show that removing either phase in LPW decreases its performance, indicating that both solution generation and code implementation phases are crucial for optimal performance. See Appendix E and F for a discussion on the accuracy of LLM-generated plans, plan verifications and refined programs in LPW.
@@
5 Cost-Performance Analysis
Figure 5 compares Pass@1 accuracy against the average token cost per program for LDB and LPW across five benchmarks using GPT-4o. When analyzing the cost for LDB, we include the tokens used to generate the seed programs, which account for about 2% of its total token consumption. LDB consumes fewer tokens per problem but achieves lower accuracy. When measured by the accuracy-cost ratio, computed as Pass@1 accuracy divided by the total tokens used, LDB shows better ratios on HumanEval and MBPP benchmarks. On LiveCode, APPS, and CodeContests benchmarks, LDB and LPW exhibit similar token usage per problem, while LPW displays notably higher accuracy. As a result, LPW realizes higher accuracy-cost ratios of 0.60% on LiveCode, 0.43% on APPS, and 0.18% on CodeContests per 1000 tokens, compared to LDB, which reports 0.50% on LiveCode, 0.39% on APPS and 0.14% on CodeContests per 1000 tokens. The lower ratios of LDB arise from insufficient refinements, where multiple ineffective debugging iterations consume significant token resources, yet the generated program remains flawed.
Figure 6 illustrates the variation in Pass@1 accuracy with token consumption across different approaches on the APPS and CodeContests benchmarks using GPT-4o. We introduce Repeated Sampling as an enhanced Baseline. For each problem, it repeatedly samples program solutions from the LLM until either the token consumption exceeds that of LPW, or a solution passes all visible tests and is validated on hidden tests. Repeated Sampling and LDB initially improve accuracy with increased tokens but show negligible improvement after around $10^{3.8}$ tokens on the APPS benchmark and $10^{3.3}$ tokens on the CodeContests benchmark. In contrast, LPW starts with high token consumption for initial plan and verification generation, resulting in a sharp accuracy increase that quickly surpasses Repeated Sampling and LDB after around $10^{4.6}$ tokens on the APPS benchmark and $10^{4.8}$ tokens on the CodeContests benchmark. Repeated Sampling is allocated the same token budget as LPW, while its final accuracy remains lower than LPW on both benchmarks, highlighting the benefits of plan and plan verification in generating high-quality initial code and subsequent refinements. Challenging benchmarks align with LPW usage scenarios, where the precise natural language solution described in the plan and plan verification is essential for logical consistency and understanding non-trivial bugs in the program, particularly when problems involve complex logical reasoning steps.